A salt solution of 15 l contains 20% salt,how much water you must add to salt solution in order to dilute it to 15% salt?
Answer by stanbon(75887)

A salt solution of 15 l contains 20% salt,how much water you must add to salt solution in order to dilute it to 15% salt?
-----
Equation:
water + water = water
0.80*15 + 1*x = 0.85(15+x)
80*15 + 100x = 85*15 + 85x
15x = 5*15
x = 5 liters (amt. of water to add)
-------------
Cheers,
Stan H.
--------------